Tarzan mass of 85 kg swings down from a tree limb on the end of the 20 m vine. His feet touch the ground 4.0m below the limb how fast is Tarzan moving when he reaches the ground
1
Expert's answer
2020-11-16T06:37:45-0500

m = 85 kg

h = 4 m

"E_p = m \times g \times h"

"E_p=85 \;kg \times 9.8 \;m/s^2 \times 4 \;m = 3332 \;J"

"E_p= \frac{1}{2}m \times v^2"

"3332= \frac{1}{2} \times 85 \;kg \times v^2"

"3332=42.5 \times v^2"

"\frac{3332}{42.5}=v^2"

"v^2 =78.4"

"v=8.85\; m/s"